if a triangle has sides of 6,8 and 12, what is the perimeter of the triangle formed by connecting the midpoints of the sides of the original triangle?




Answer by oscargut(2103)

Is the half of the perimeter of the original triangle
Is (6+8+12)/2 = 26/2 = 13
Answer: 13
If you need more help with math questions/assignments/tests
You can ask me more using this e-mail (mthman@gmail.com) or you can visit my site:
http://www.online-maths-help.com/
(Payment is using paypal)